A map-maker wants to represent an area of roughly 10,000ft^2 using the scale 1inch=50ft . What area will the map have?
Answer by jim_thompson5910(35256)

1inch=50ft

square both sides to get

1 square inch = 2500 square feet


So the ratio of the larger area to smaller area is 2500/1


(area in real life)/(area on map) = Ratio

10000/x = 2500/1

10000*1 = 2500x

10000 = 2500x

2500x = 10000

x = 10000/2500

x = 4

So the area of the map will be 4 square inches